OpenStudy (mikee):

A 12lb shankless ham contains 16 servings. What is the rate in servings per pound?

OpenStudy (anonymous):

12/16 = 6/8 = 3 / 4 servings per pound

OpenStudy (anonymous):

oops 16/12 = 8 / 6 = 4 / 3 servings per pound

OpenStudy (mikee):

Thank you so the answer is 4/3 servings per pound?

OpenStudy (anonymous):

16/12 = 8 / 6 = 4 / 3 servings per pound